# Quantum Computing in Wireless Communications and Networking: A Tutorial-cum-Survey

### v1

@@ II-E QC in Noisy Intermediate-scale Quantum Era
As the name suggests, the mixing Hamiltonian $\hat{B}$ offers “mixing” operations. If regarding $\beta$ as a time-dependent parameter, $e^{-i\beta\hat{B}}$ actually describes a continuous quantum walk on the graph $\hat{B}$, on which the walker can reach any connected nodes from the initial state (node). In other words, the mixing operation is a specific unitary operation evolving the initial state to a superposition state mixed with other feasible states. In the scheme for unconstrained problems, or the QUBO-based scheme for constrained problems, the mixing Hamiltonian is always defined as $\sum_{i}\sigma^{x}_{i}$. In physics, this operator is called the transverse field operator, and its corresponding graph is a hypergraph, which connects pairwise nodes by $Hamming\ distance=1$.
In the alternative scheme for tackling constrained problems, mixing Hamiltonian designing is more subtly, since eliminating penalty terms and restricting the dynamics to the feasible space can get potential benefits of faster convergence with lower cost. Hadfield et al. proposed the seminal approach – Quantum Alternating Operator Ansatz, in which they constructed XY-mixer, Partial mixer, etc., for varied constraint types [64]. In recent two years, there are emerging many variants mixing Hamiltonian [65], which aims at evolving the initial state to the target state through smaller $p$ iterations via the mixing ability. A smaller $p$ means a shallower quantum circuit, thereby more efficiency and less suffering from decoherence.
Problem Hamiltonian $\hat{C}$ is used to encode the objective function, i.e. $\bra{x}\hat{C}\ket{x}=C(x)$, where $x$ is a bit string whose components “0/1” indicates the corresponding combinatorial variable is included in the solution or not; $C(x)\in\mathbb{R}$ is the objective function under the variable $x$. Depending on the optimization problem at hand, the highest/lowest energy state |$\ket{x^{*}}$ of Ising Hamiltonian $\hat{C}$ encodes the optimal solution. For Ansatz-based schemes, one needs to formulate each component of $x$ with $\frac{I-\sigma_{i}^{z}}{2}$ (the so-called spin to binary relation) and then use it to generate matrix $\hat{C}$; for QUBO-based schemes, still needs to construct extra penalty terms in $\hat{C}$, the tricks may refer to [66].
@@
In terms of $\hat{U}_{B}$ and $\hat{U}_{C}$, one translates these operators to circuits for execution. The initial state goes through the circuit, gets the final state |$\ket{x}$, does measurement $\bra{x}\hat{C}\ket{x}$, and then calibrates the values of $\vec{\beta}$ and $\vec{\gamma}$ for the next cycle until |$\ket{x}$ converges to |$\ket{x^{*}}$. This procedure, like VQE, also involves the choice of classical strategies and error migration methods, which we will discuss in the subsequent sections.
Besides, note that implementing a QAOA-based algorithm in a real QC device is more challenging since the operator form of its variant versions is more complex. Moreover, current QC devices differ widely in qubits-connecting topology, native gates set, fidelity, etc., and compiling software offered by vendors is far from being able to abstract these underlying physical characteristics. Thus, a lot of additional hard work needs to be done to achieve good performance [35].
3) Quantum Machine Learning
Quantum machine learning, with which physicists have been fascinated for more than two decades [67], holds a solid foundation since the matrix manipulating techniques used in AI are well-studied in quantum theory. For this reason, and with the AI wave, people hope to further fuel the surge in AI research through the rich theory and technique of quantum technology. And this is indeed the case, many remarkable achievements have emerged in recent years [68]. For example, QSVM has a logarithmic runtime in both training and classification stages, achieving an exponential acceleration compared with the classical counterpart [42]. Nevertheless, the current capabilities of QC devices are lagging behind since its core subroutine – quantum principal component analysis (QPCA) [69] employs quantum Fourier transform, in which the controlled unitary gate is decomposed into a deep quantum circuit far beyond the capacity of QC devices in NISQ era. Given there is a long way toward large-scale fault-tolerant QC, quantum machine learning and other quantum algorithms require finding ways to adapt to the limited scale of imperfect QC resources [70]. That means the quantum algorithms running on the shallow-depth quantum circuit could still achieve some practical benefit. To this end, the hybrid workflow – uses parameterized quantum circuits (PQC) to run on a quantum computer and then optimizes parameters on a classical computer, becomes a typical QNN architecture. To some extent, PQC is the synonym of QNNs, and VQE and QAOA could be regarded as two special cases of PQC [71]. The quantum machine learning literature is so rich that our discussion is not going to cover all research branches. Instead, we focus on some hot topics: classification, generative AI, and RL.
@@
Classification, broadly also including clustering, is the most studied topic in machine learning. In PQC-based frameworks, quantum subroutines replace one or several steps in a classical neural network aims at improving the neural network’s abilities of expressibility, trainability, generalization, etc. [71], which Llyod calls quantum embedding [72]. Fig. 9 shows a typical architecture of such frameworks. The leftmost block denotes a preprocessing step on the classical computer to extract feature data. This step can be simple as, for example, in terms of the gray value of an image pixel, embedding it as the rotation angle of the rotation gate in eq. (45) for the subsequent quantum processing [73], where $a_{i}\in[0,255]$ is the gray value of a pixel..
@@
Still, it can become more complex. [74, 75] employ quantum autoencoders to extract distinguishing characteristics from quantum states in a classical latent space, so as to use reduced quantum data for clustering and classification. Quantum embedded circuits for encoding is usually neither optimized nor trained, which would affect the computational power of QNNs [76].
Figure 9: The schematic diagram of typical hybrid quantum convolutional neural network architectures.
The abstract notation $U(\phi)$ and $U(\theta)$ in the middle quantum processing block represent two unitary operations for mapping classical data to the quantum domain and generating entanglement, respectively. $U(\phi)$ and $U(\theta)$ could be varied a lot in terms of different scenarios. In quantum convolutional neural networks (QCNNs), $U(\phi)$ and $U(\theta)$ are combined together as a quantum convolution layer implemented by a cascade of several layers of $R_{y}(\theta)$ rotation gate and CNOT gates [73]. Another version of QCNNs combines the multi-scale entanglement renormalization ansatz (MERA) to implement a pooling layer besides the quantum convolutional layer [77]. One point we would like to emphasize is that entanglement is essential to achieve a quantum advantage using hybrid neural networks. In a quantum transformer architecture, the entanglement is generated by RBS gates, which can be decomposed further by four $Hadamard$ gates, two $R_{y}$ rotation gates, and two two-qubit $CZ$ gates [78, 73].
After quantum data are remeasured, they become classical data again to continue the remaining operations in classical neural networks. Depending on the problem at hand, varied objective/loss functions are computed, and the result is used to calibrate the variational values $\boldsymbol{\phi}$ and $\boldsymbol{\theta}$ until convergence. This process involves training iteratively and dealing with noise and gate errors that all PQC-based algorithms confront. Therefore, we will discuss this in a subsequent independent section.
@@
In recent years, there is no doubt that the hottest AI technology is generative AI, which generates new samples by taking the underlying probability distribution of a dataset such as images, videos, etc. From the Boltzmann machine [79], to the latest Generative Adversarial Networks (GAN) [80] and Denoising Diffusion Probabilistic Model (DDPM) [81], all have corresponding quantum versions.
The quantum generative learning models (QGLMs) seem to be more versatile than their classical counterparts. Besides image generation [82], there are works discuss quantum state estimation [83], continuous distributions generation [84], and distribution learning [79] etc., show their ability for accomplishing what the classical tools are incapable of.
QGLMs adopt the same workflow as their classical counterparts. For example, the Quantum QGLMs (QuDDPM) [81], shown in Fig. 10, is also composed of forward noisy and backward denoising process. The forward noisy diffusion process exploits a quantum scrambling circuit that evolves the sample data towards a random ensemble of pure states by applying a series of random unitary gates independently. On the other hand, the backward denoising process reduces noise gradually towards the final generated data by performing projective measurements after applying unitaries on ancilla qubits. This process ensures that the generated data maintains purity and mimics the sample data. The integration of QC techniques and diffusion models offers benefits that are not only suitable for efficient training and learning of quantum data but also improve performance in classical tasks. As an example, [85] show two quantum diffusion models – Q-Dense and QU-Net outperform classical counterparts in terms of performance metrics like FID, SSIM, and PSNR.
@@
The key challenge in QGLMs is exploring their learnability in terms of generalization, expressivity, and trainability. The generalization and expressivity of QGLMs are less studied compared to quantum discriminative learning models. The definition of generalization in quantum generative learning is vague under different model settings, making it difficult to study the generalization of QGLMs. The trainability of QGLMs is also challenging due to the non-convex loss landscape and the presence of barren plateaus. The performance of QGLMs can be significantly affected by quantum hardware imperfections such as gate noise and measurement error. When the level of noise exceeds a threshold, the power of near-term quantum computers may be inferior to classical computers. Quantum system noise can also lead to diverged optimization in QGLMs.
@@
Given that QC can enhance computational power, thus naturally, combining QC and RL was conceived and has been studied for more than a decade [86]. QRL can be roughly divided into three categories according to the proportion of ‘quantum’: the first is so-called inspired RL, which basically belongs to the category of classical heuristics. The second is PQC-based RL, which can run on NISQ devices and will be the focus of our subsequent discussion. The last one is full-quantum RL. As the name suggests, algorithms in this category only work well on fault-tolerated quantum computers.
PQC-based RL can be further classified in terms of the role of PQC in the pipeline, one is the value-function approximation approaches; one is policy approximation approaches, and the combined approximation for value and policy, i.e., actor-critic (AC) approaches. A typical PQC-based QRL pipeline is summarized in Fig. 11. The hybrid agent observes the initial environment and applies some pre-processing to encode the observation into the feature map $U(\phi)$. Through a $\theta$-parameterized unitary operation $U(\theta)$, the current state |$\ket{s_{t}}=U(\theta)U(\phi)\ket{0}^{\otimes n}$ can be gotten. Then depending on the measurements of the action-selection observable $\bra{s_{t}}O_{a}\ket{s_{t}}$, the state-action value function $Q_{\theta}(s_{t},a_{t})$ or the policy $\pi_{\theta}(a_{t}|s_{t})$ will navigate the next favorable action $a_{t+1}$. Finally, the cumulative reward $r_{t}$ computed by interacting with the environment is iterated into a classical optimizer for consecutive parameter updates and optimization until convergence. The overall iterative pipeline is similar to the classical DRL, except that state change occurs in the quantum domain, where the variational unitaries could be regarded as a kind of QNNs for approximation.
@@
Applying quantum approaches to RL presents several challenges. One key challenge is the limited exploration of the true potential of near-term quantum approaches in RL, as existing works have struggled to solve classical benchmarking tasks using PQCs. Another challenge is achieving sufficient performance in RL environments when using PQCs as value-function or policy approximators, as demonstrated in recent works. Additionally, the complexity and resource-intensive nature of quantum computations pose challenges in terms of scalability and practical implementation in RL tasks. Furthermore, the interpretability of quantum learning models is a concern, as the quantum nature of computations make these models difficult to understand compared to classical ones. Lastly, the accessibility of quantum computers for training QRL models potentially exacerbates existing inequalities in AI development and usage.
@@
• Barren plateaus: Barren plateaus (BP) are characterized by an exponential vanishing of gradients across the optimization landscape. These plateaus not only impede gradient-based optimizers but also pose challenges for gradient-free algorithms like Nelder-Mead and COBYLA, as the flat landscape tends to exponentially dampen the cost function [87]. The emergence of BP limits the potential of PQCs to achieve quantum advantages. This is because the exponential decline of gradients necessitates exponential resources for optimization, whereas quantum algorithms typically aim for polynomial scaling of the system. Consequently, as problem sizes increase, PQCs become increasingly difficult to train.
@@
BP occurs in varied cases: random initialization in parametric circuit [88], objective function defined in terms of global observables [89], entanglement-induced [90], noise-induced [91], and PQC structure-induced [92], etc. For various scenarios, people have come up with methods to alleviate or avoid the occurrence of BP. Layerwise learning, transfer learning, and Bayesian learning have been used to find a favorable initial set of parameters for mitigating BPs [93]. [94] proposed an algorithm called generalized quantum assisted simulator, which does not require a classical-quantum feedback loop, to bypass the BP problem. [95] established a fundamental connection between expressibility and trainability by extending the concept of barren plateaus to arbitrary ansatz, which provides insights into designing strategies for problem-inspired ansatz to avoid BPs. [96] show that over-parametrization in QNNs can mitigate BPs.
@@
Note that although there has been a wealth of research on the BP problem, little is known about the optimization landscape of a generalized PQC-based algorithm. To the best of our knowledge, the only one that can be sure to avoid BP is the work [97]. The paper rigorously analyzes the gradient scaling for parameters in QCNNs architecture. It finds that the variance of the gradient vanishes no faster than polynomials, indicating that QCNNs do not exhibit barren plateaus. This result is significant as it provides an analytical guarantee for the trainability of randomly initialized QCNNs.
• Run quantum algorithms in real quantum computers: As the first step, the programmer must decompose unitary operators (matrix) in the quantum algorithm into small ones – single-qubit and double-qubit gates. For a generalized unitary matrix $e^{-iH}$, till now, we only know that Hamiltonian $H$ must be sparse, or low-rank [69], then it can have a polynomial decomposition. For decades, numerous related works have struggled to find methods for more efficient decomposition. Next, the programmer translates logical gates into native gates that the specific hardware supports. Since the underlying supports for current QC vary widely in technology (superconducting, ion trap, optical photon, etc.), and even with the same technology, they still differ in low level, a lot of extra work needs to be done for the optimization [98]. For example, IBM and Rigetti ’s machines are both grounded in superconducting technology. The former has the native CNOT gate, while the latter needs to translate $CNOT(A,B)$ gate into a sequence of rotations and CZ operations: $R_{z}(\pi/2)B;R_{x}(\pi/2)B;R_{z}(\pi/2)B;CZ(A,B);R_{z}(\pi/2)B;\\ R_{x}(\pi/2)B;R_{z}(\pi/2)B$ [99]. Compilers provided by vendors, such as IBM’s Qiskit, can translate multi-control gates into native gates customized for their architecture.
@@
Finally, programmers need to find ways to mitigate various types of errors. In the NISQ era, hardware resources cannot support the significant consumption of applying quantum error correction code. Quantum error mitigation, which suppresses errors via classical postprocessing techniques, is the only practical approach to handle errors at this stage[70]. It requires the noise model in advance and then estimates parameters for the expected value of the noise-free model by analyzing how the noise model behaves under different noise levels. The other is probabilistic error cancellation based on gate set tomography [100], which does not know the precise physical noise model in advance. It combines linearly independent basis set operations and gate set tomography to fully remove the impact of localized Markovian errors by systematically measuring the effect of errors. This technique has been successfully implemented in trapped-ion systems and superconducting systems.

### v3

@@ II-E QC in Noisy Intermediate-scale Quantum Era
The mixing Hamiltonian $\hat{B}$ is designed to perform ”mixing” operations. When $\beta$ is treated as a time-dependent parameter, $e^{-i\beta\hat{B}}$ describes a continuous quantum walk on the graph defined by $\hat{B}$. This allows the quantum state to evolve from the initial state to a superposition of other feasible states. In simpler terms, the mixing operation represents a specific unitary transformation that drives the system from its initial configuration into a mixed superposition of possible solutions. In the case of unconstrained problems, or QUBO-based constrained problems, the mixing Hamiltonian is typically defined as $\sum_{i}\sigma^{x}_{i}$. In physics, this operator is referred to as the transverse field operator, and its corresponding graph is a hyper-graph where nodes are connected pairwise by a Hamming distance of 1.
For constrained optimization problems, designing the mixing Hamiltonian becomes more intricate. By eliminating penalty terms and restricting the quantum evolution to the feasible subspace, one can achieve faster convergence at a lower computational cost. QAOA introduces specialized mixers such as XY-mixer and partial mixers for different types of constraints [55]. Recent works have introduced further variants of the mixing Hamiltonian [56], focusing on reducing the required number of iterations $p$ through enhanced mixing capabilities. A smaller $p$ corresponds to a shallower quantum circuit, which is beneficial for maintaining quantum coherence and reducing gate errors.
Problem Hamiltonian $\hat{C}$ encodes an objective function to be optimized, where $\bra{x}\hat{C}\ket{x}=C(x)$, and $x$ is a bit string that indicates the inclusion of combinatorial variables in the solution. $C(x)\in\mathbb{R}$ is the value of the objective function for a given bit string $x$. Depending on specific optimization problems, the highest or lowest energy eigenstate |$\ket{x^{*}}$ of the Ising Hamiltonian $\hat{C}$ represents the optimal solution. For Ansatz-based schemes, each component of $x$ is formulated using $\frac{I-\sigma_{i}^{z}}{2}$, known as the spin-to-binary conversion, which is used to generate matrix representation of $\hat{C}$. In QUBO-based schemes, additional penalty terms have to be constructed within $\hat{C}$ [57].
@@
In terms of the unitary operators $\hat{U}_{B}$ and $\hat{U}_{C}$, they are translated into quantum circuits for execution on a quantum computer. The initial state is prepared and evolved through the circuit, producing a final state |$\ket{x}$. A measurement is performed to evaluate $\bra{x}\hat{C}\ket{x}$, after which parameters $\vec{\beta}$ and $\vec{\gamma}$ are adjusted for the next iteration. This process continues until final state |$\ket{x}$ converges to the optimal solution |$\ket{x^{*}}$. Similar to VQE, this procedure involves the selection of classical optimization strategies and error mitigation techniques, which will be discussed in subsequent sections.
It is important to note that implementing QAOA-based algorithms on real quantum hardware presents additional challenges. The operator forms of different QAOA variants are often complex, making the quantum circuit deeper and more prone to errors. Furthermore, current quantum devices vary significantly in qubit connectivity topology, native gate sets, and operational fidelity. Compiling software provided by vendors is not yet capable of abstracting these hardware-specific characteristics efficiently.
3) QML
QML, a fascinating cross-research field, holds a solid foundation since the matrix manipulating techniques used in AI are well-studied in quantum theory. Over the past two decades, research in this area has achieved significant milestones. For example, QSVM exhibits a logarithmic runtime in both training and classification phases, offering an exponential speedup over its classical counterpart [36]. However, this advancement remains theoretical, since a key subroutine of QSVM, quantum principal component analysis [58], relies on the quantum Fourier transform, where the decomposition of controlled unitary gates results in quantum circuits too deep for near-term devices in the NISQ era. Thus, QML, along with other quantum algorithms, must adapt to these constraints and operate efficiently within the limited capabilities of present QC resources. This adaptation has led to the hybrid workflow model, where PQC runs on quantum computers while classical computers optimize the parameters. PQC has effectively become synonymous with QNNs, and algorithms like VQE and QAOA can be considered special cases of PQC [59]. Given extensive QML literature, this discussion focuses on a few key areas, particularly classification, generative AI, and RL.
@@
Classification, including clustering, is one of the most extensive studies in ML. In PQC-based frameworks, quantum subroutines replace one or more steps in a classical neural network, with the aim of enhancing the network expressibility, trainability, and generalization capabilities [59, 60]. Fig. 9 illustrates a typical architecture of such frameworks. The leftmost block represents a pre-processing step, say feature extraction, on a classical computer. This step may be straightforward. For example, in the case of an image, the gray value of each pixel is embedded as the rotation angle of a quantum rotation gate in eq. (48) for subsequent quantum processing [61]. Here, $a_{i}\in[0,255]$ represents the gray value of a pixel:
@@
The process could be more complex. For instance, [62] utilizes quantum auto-encoders to extract key features from quantum states into a classical latent space. This enables clustering and classification using reduced quantum data.
Figure 9: The schematic diagram of typical hybrid quantum convolutional neural network architecture.
The abstract notations $U(\phi)$ and $U(\theta)$ in the middle quantum processing block denote unitary operations for mapping classical data to the quantum domain and generating entanglement, respectively. The specific forms of $U(\phi)$ and $U(\theta)$ can vary depending on scenarios. In quantum convolutional neural networks (QCNNs), $U(\phi)$ and $U(\theta)$ are combined to create a quantum convolution layer, typically realized by a series of rotation gate $R_{y}(\theta)$ and CNOT. Another variant of QCNNs incorporates multi-scale entanglement renormalization ansatz to implement a pooling layer in addition to the quantum convolutional layer. It is important to highlight that entanglement plays a crucial role in achieving a quantum advantage with hybrid neural networks. In quantum transformer architectures, entanglement is generated using gate RBS, which can be further decomposed into four Hadamard gates, two rotation gates $R_{y}$, and two two-qubit gates $CZ$ [63].
After quantum data are remeasured, they are converted back into classical data to proceed with the remaining operations in classical neural networks. Depending on specific problems, various objective or loss functions are computed. The results are then used to adjust the variational parameters $\boldsymbol{\phi}$ and $\boldsymbol{\theta}$ until convergence is achieved. This iterative training process also involves addressing noise and gate errors, challenges that are common to all PQC-based algorithms. A detailed discussion on these aspects is provided in a subsequent section.
@@
In recent years, generative AI has emerged as a leading technology, capable of generating new samples by modeling the underlying probability distribution of datasets such as images and videos. This includes a range of methods from the Boltzmann machine [64] to contemporary approaches like generative adversarial networks (GANs) [65] and denoising diffusion probabilistic models (DDPM) [66]. Each of these techniques has its quantum counterpart.
Quantum generative learning models (QGLMs) have demonstrated greater versatility compared to classical models. They have been applied to quantum state estimation [67], continuous distribution generation [68], and distribution learning [64], showing their capability to tackle problems that classical methods struggle with.
QGLMs generally follow the same workflow as their classical equivalents. For instance, the quantum version of the denoising diffusion probabilistic model (QuDDPM) [66], depicted in Fig. 10, consists of both forward noisy and backward denoising processes. In the forward noisy diffusion process, a quantum scrambling circuit is utilized to evolve sample data towards a random ensemble of pure states. It is achieved by applying a series of random unitary gates independently. Conversely, the backward denoising process involves gradually reducing noise to obtain the final generated data. This is done by performing projective measurements after applying unitaries on ancilla qubits. This approach ensures that the generated data remains pure and closely resembles the sample data. The integration of QC techniques with diffusion models provides significant advantages, not only for efficient training and learning of quantum data but also for enhancing performance in classical tasks. For example, [69] demonstrates that two quantum diffusion models, Q-Dense and QU-Net, outperform their classical counterparts.
@@
The primary challenge in QGLMs is assessing their learnability, particularly in terms of generalization, expressivity, and trainability. Compared to quantum discriminative learning models, the generalization and expressivity of QGLMs have been less extensively studied. The concept of generalization in the context of quantum generative learning remains ambiguous across different model settings. Trainability is another significant challenge for QGLMs, due to in part the non-convex nature of loss landscapes and the presence of barren plateaus, where gradients become small and hinder effective training. Furthermore, the performance of QGLMs is heavily influenced by quantum hardware imperfections, including gate noise and measurement errors. When noise levels exceed a certain threshold, near-term quantum computers may perform worse than classical counterparts. Additionally, quantum system noise can lead to divergent optimization processes in QGLMs, further complicating their practical implementation.
@@
Given that QC can enhance computational power, combining QC and RL has been studied for more than a decade. QRL can be roughly divided into three categories according to the proportion of ‘quantum’: the first is so-called inspired RL, which basically belongs to the category of classical heuristics. The second is PQC-based RL, which can run on NISQ devices and is the focus of our subsequent discussion. The last one is full-quantum RL. As the name suggests, algorithms in this category only work well on fault-tolerated quantum computers.
PQC-based RL can be further classified in terms of the role of PQC in the pipeline, one is the value-function approximation approaches; one is policy approximation approaches, and the combined approximation for value and policy, i.e., actor-critic (AC) approaches. A typical PQC-based QRL pipeline is summarized in Fig. 11. The hybrid agent observes the initial environment and applies pre-processing to encode the observation into the feature map $U(\phi)$. A $\theta$-parameterized unitary operation $U(\theta)$ then generates the current state |$\ket{s_{t}}=U(\theta)U(\phi)\ket{0}^{\otimes n}$. Based on the measurements of the action-selection observable $\bra{s_{t}}O_{a}\ket{s_{t}}$, either the state-action value function $Q_{\theta}(s_{t},a_{t})$ or the policy $\pi_{\theta}(a_{t}|s_{t})$ determines the next favorable action $a_{t+1}$. The cumulative reward $r_{t}$ obtained from interacting with the environment is subsequently fed into a classical optimizer for iterative parameter updates and optimization until convergence. This iterative pipeline resembles classical DRL, with the distinction that state changes occur in the quantum domain, where the variant can be viewed as a form of QNNs for approximation.
@@
Applying quantum approaches to RL presents several challenges. One key challenge is the limited exploration of the true potential of near-term quantum methods in RL, as existing research has struggled to solve classical benchmarking tasks using PQC. Additionally, achieving sufficient performance in RL environments with PQCs as value-function or policy approximators has proven difficult. The complexity and resource-intensive nature of quantum computations further complicate their scalability and practical implementation in RL tasks. Quantum computations require substantial resources, which can hinder their feasibility for real-world applications. Moreover, the interpretability of quantum learning models is a concern. The quantum nature of these computations makes them difficult to understand compared to classical models, complicating efforts to improve models. Finally, the accessibility of quantum computers for training QRL models could exacerbate existing inequalities in AI development and usage, potentially creating disparities in the ability to leverage quantum-enhanced technologies.
@@
PQC-based algorithms, or hybrid variational quantum algorithms, typically involve a classical optimization subroutine to train parametrized quantum circuits and identify the optimal values for the objective function. A significant challenge in this context is the presence of barren plateaus, which can severely limit training efficiency. In the following section, we review related work addressing the conditions under which barren plateaus arise and explore strategies to avoid or mitigate them. Additionally, the practical implementation of quantum algorithms on NISQ devices presents its own set of challenges. Due to variations in QC infrastructure and the presence of various types of noise, executing quantum algorithms on actual hardware is substantially more complex than running them on simulators. This challenge is discussed in detail in the subsequent section.
• Barren plateaus:
@@
Barren plateaus (BP) are characterized by an exponential vanishing of gradients across the optimization landscape. These plateaus not only impede gradient-based optimizers but also pose challenges for gradient-free algorithms like Nelder-Mead and COBYLA, as the flat landscape tends to exponentially dampen the cost function [70]. The emergence of BP limits the potential of PQCs to achieve quantum advantages. This is because the exponential decline of gradients necessitates exponential resources for optimization, whereas quantum algorithms typically aim for polynomial scaling of the system.
@@
BP occurs in varied cases [71] like random initialization in parametric circuit, objective function defined in terms of global observables, entanglement-induced, noise-induced, and PQC structure-induced. Layerwise learning, transfer learning, and Bayesian learning have been used to find a favorable initial set of parameters for mitigating BP [72]. [73] proposes an algorithm called generalized quantum assisted simulator, which does not require a classical-quantum feedback loop. [74] establishes a fundamental connection between expressibility and trainability by extending the concept of BP to arbitrary ansatz, which provides insights into designing strategies for problem-inspired ansatz to avoid BP. [75] show that over-parametrization in QNNs can mitigate BP.
@@
Note that although there has been a wealth of research on the BP problem, little is known about the optimization landscape of a generalized PQC-based algorithm. To the best of our knowledge, the only one that can be sure to avoid BP is the work [76]. The paper rigorously analyzes gradient scaling for parameters in QCNN architecture. It finds that the variance of the gradient vanishes no faster than polynomials, indicating that QCNNs do not exhibit BP. This result is significant as it provides an analytical guarantee for the trainability of randomly initialized QCNNs.
• Run quantum algorithms in real quantum computers:
The first step in implementing a quantum algorithm involves decomposing unitary operators (matrices) into smaller components—typically single-qubit and two-qubit gates. For a generalized unitary matrix $e^{-iH}$, it is currently known that the Hamiltonian $H$ must be sparse or low-rank to allow for a polynomial decomposition [58]. Over the past few decades, significant efforts have been dedicated to discovering more efficient methods for such decompositions. Once the unitary operators are decomposed, the next step is translating logical gates into native gates supported by the specific quantum hardware. Since QC platforms vary widely in underlying technologies (superconducting qubits, ion traps, optical photons, etc.), and even within the same technology there are differences at lower levels, considerable effort is required for optimization [77]. For instance, while both IBM and Rigetti machines are based on superconducting technology, their native gate sets differ. IBM natively supports the CNOT gate, whereas Rigetti requires the translation of a $CNOT(A,B)$ gate into a series of rotation and CZ operations. Vendor-provided compilers, such as IBM’s Qiskit, are capable of translating multi-controlled gates into native gates that are optimized for specific hardware architectures.
@@
Finally, programmers must address various types of errors inherent in quantum systems. In the NISQ era, quantum hardware lacks the resources to fully support intensive demands of quantum error correction codes. Thus, quantum error mitigation, which reduces errors through classical postprocessing techniques, is currently the most practical solution for handling errors in NISQ devices [78]. One common method requires knowledge of the noise model beforehand. It estimates parameters for the expected value of the noise-free model by analyzing the behavior of the noise model under different noise levels. Another technique is probabilistic error cancellation, which does not rely on precise knowledge of physical noise models in advance. It is based on gate set tomography and uses a linearly independent basis of operations to systematically remove the effects of localized Markovian errors [79]. By measuring the impact of these errors, probabilistic error cancellation has been successfully implemented in both trapped-ion and superconducting systems.